7Binomial
. Find . Give the answer to 4 decimal places.
At least one success is easier found as 1 minus the chance of zero successes. Here P(X=0) = 0.8⁴ = 0.4096, so P(X≥1) = 1 − 0.4096 = 0.5904.
. Find . Give the answer to 4 decimal places.
At least one success is easier found as 1 minus the chance of zero successes. Here P(X=0) = 0.8⁴ = 0.4096, so P(X≥1) = 1 − 0.4096 = 0.5904.